Use the graph of f to sketch each graph.

(0,5) (-3,0) (3, 0), -5 5 (6,-4) (6, -4) ()yfx5) 1아 10 A X -5 -10 5 10 -10 10 -10 -10 1아 X -10 5 5 10 -10 -5 5 10 -10 -10-f(x) (b) 5 y V 10 10 -10 10 -10 5 10 -10 10 10 X -10 5 10 -10 5 10 -5 -10! -1아(c) y 15k 15 1아 10 X X -15 -10 -5 10 15 -15 -10 10 15 -10 -10 -15 -15 y 15 15 10 5 X 15 -5 5 -15 -10 10 -15 -10 -5 5 10 15 -1(d) y -fx 4) y 1아 10 X 5 -5 -10 10 -10 5 10 -10 -10 1아 10 X -10 -5 -10 5 10 5 10 -10 -10(e) yf-x) y 10 10 -10 5 10 -10 5 10 -10 -10 10H 10 -5 -10 10 -10 5 10 -10 -10y fx) (f) 10 = y 151 10 10 x 15 5 -5 -15 -10 5 10 -15 -10 5 10 15 -10 -10 -15 -15 y 15H 15 10 10 A x 5 10 -15 -10 5 15 -15 -1(g) (4 1아 10 X -5 -10 10 -10 5 10 -10 -10 10 10 -5 -5 -10 5 10 -10 10 -5 -10 -10

(0,5) (-3,0) (3, 0), -5 5 (6,-4) (6, -4) ()yfx5) 1아 10 A X -5 -10 5 10 -10 10 -10 -10 1아 X -10 5 5 10 -10 -5 5 10 -10 -10
-f(x) (b) 5 y V 10 10 -10 10 -10 5 10 -10 10 10 X -10 5 10 -10 5 10 -5 -10! -1아
(c) y 15k 15 1아 10 X X -15 -10 -5 10 15 -15 -10 10 15 -10 -10 -15 -15 y 15 15 10 5 X 15 -5 5 -15 -10 10 -15 -10 -5 5 10 15 -1아 -15 -15
(d) y -fx 4) y 1아 10 X 5 -5 -10 10 -10 5 10 -10 -10 1아 10 X -10 -5 -10 5 10 5 10 -10 -10
(e) yf-x) y 10 10 -10 5 10 -10 5 10 -10 -10 10H 10 -5 -10 10 -10 5 10 -10 -10
y fx) (f) 10 = y 151 10 10 x 15 5 -5 -15 -10 5 10 -15 -10 5 10 15 -10 -10 -15 -15 y 15H 15 10 10 A x 5 10 -15 -10 5 15 -15 -10 -5 5 10 15 10b -10 -15F -15
(g) (4 1아 10 X -5 -10 10 -10 5 10 -10 -10 10 10 -5 -5 -10 5 10 -10 10 -5 -10 -10
Answer #1

WELCOME

a.

y = f(x-5)

The graph will shift 5 unit right

'C' is the right answer

b.

y = -f(x) + 5

after reversing graph will move 5 unit upwart

'C' is the right answer

c.

y = (1/2) f(x)

'B' is the right answer

d.

y = -f(x+4)

after reversing it wil move 4 unit left

'C' is the right answer